A thin 30cm candle takes 3h to burn from one end to the other. A thicker candle with the same length takes 5h to burn from one end to the other. If the thicker candle is lit at 7pm, and the thin candle is lit 30 min later, at what time will the candles be at the same height?
Answer by mananth(16946)

Thin candle burning rate is 30/3 = 10 cm/h
Thick candle burning rate = 30/5 = 6 cm/h
Thin candle burns 10-6 = 4 cm/h faster than thick candle
Thicker candle is lit at 7
thinner candle is lit at 7.30
so difference = 1/2 hour.
in 1/2 hour the thin candle would have burnt 10*1/2 = 5 cm
time both will have same height = difference in length when second candle is lighted / difference between burning rates
5/4 = 1.25 hours
m.ananth@hotmail.ca